OpenStudy (anonymous):

write an explicit formula for the arithmetic sequence 62, 59. 56, 53

OpenStudy (anonymous):

nth term =a+(n-1)*d where a is the first term here its 62 and d is the common difference here u can calculate d by subtracting the next number in sequence with its preceding number here d is 59-62 = -3 , also 56 - 59 = -3 its common , thats y we call it common difference :) now nth term = 62+(n-1)*(-3) = 62 -3n+3 = 65-3n

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

My pleasure and thanks for the medal